# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 1.2.2. Boundary HEMs
Following the same mechanism as in the bulk case, Theorem 1.3 should lead to PDEs for conformal blocks on arbitrary Riemann surfaces with boundaries and punctures. Since the theory is not as well developed for boundary LCFT, we postpone this discussion to future work. Let us just mention that these PDEs specialise to the result of [Ang23] when: (i) the cosmological constants are evaluated on the FZZ conic section, (ii) the Riemann surface is the disc with an arbitrary number of punctures on the disc or the boundary, (iii) conformal blocks are replaced by correlation functions. Therefore, the equations we get are a three-fold generalisation of [Ang23]. The proof of [Ang23] is drastically different from ours since it uses the mating-of-trees theory [DMS21]: it relies on the clever observation that the BPZ differential operator is the generator of a certain Schramm-Loewner evolution. Using this observation, the BPZ equation is a consequence of the fact that the correlation function evolves as a martingale under the dynamics generated by this operator. In this context, the quantities $\cos(\pi\frac{\gamma^{2}}{4})$ and $\sin(\pi\frac{\gamma^{2}}{4})$ appearing in the RHS of the $(2,1)$-HEM have a special meaning: they are respectively the correlation and the speed of the Brownian motions corresponding to the evolution of the two quantum lengths (parametrised by quantum area). In our proof, these quantities are related to residues of certain Selberg integrals on the degenerate weights (Lemma 3.5).
Finally, we are aware of a work in progress by Cerclé [Cer] (private communication with the author) which gives a new proof of Ang’s result using the usual framework of Ward identities. Cerclé introduces a “desendant field at level 2” which can be thought of as the descendant state $\Phi_{\alpha}(\mathbf{S}^{0}_{2Q-\alpha}\mathds{1})$ from this paper. The insertion of this state combined with Ward identities leads to the BPZ equation. In our approach, the Ward identities will be replaced with the conformal blocks machinery that will be developed thanks to the boundary version of the conformal bootstrap. Indeed, this machinery gives a way to streamline the Ward identities without computations and independently of the geometric setup, using the framework of Virasoro uniformisation.
1.3. Outline
The strategy for the proof of Theorems 1.1 and 1.3 is the following
@@ 1.3. Outline
    We express the singular state as certain singular integral of the correlation function (Sections 2.2 and 3.2). These expressions are reminiscent to those found in [BW23], namely they look like primary fields with additional $\gamma$-insertions integrated over the disc (see (2.1) and (3.3)).
@@
    Contrary to [BW23], these integrals have a simple pole at $\alpha_{1,2}$, $\alpha_{2,1}$. The presence of this pole is the crux of the HEM: the RHS of the HEMs are given by the residues of these integrals. To compute this residue, we rely on fusion estimates for correlation functions (Propositions 2.6 and 3.6), and exact formulas for Dotsenko-Fateev/Selberg integrals.
@@
    The general theory tells us that the Poisson operator is analytic away from the Kac table. Since $\alpha_{1,2}<\alpha_{2,1}$, the pole at $\alpha_{1,2}$ is easier to study. To study the pole at $\alpha_{2,1}$, one needs to find a probabilistic expression for the analytic continuation beyond the pole at $\alpha_{1,2}$. This done in Propositions 2.7 and 3.7, using a method similar to the one in [BW23, Section 3.3].
@@
The RHS of (1.3) has a limit as $\gamma\to\sqrt{2}$ from below, but this limit is deceptively simple: the constant prefactor has a pole (as a function of $\gamma$), which is compensated by the zero of the primary state. Thus, the limit involves the derivative state $\Psi_{Q}^{\prime}=\underset{\alpha\nearrow Q}{\lim}\,\frac{\Psi_{\alpha}}{\alpha-Q}$. It is easy to see that this state has the probabilistic formula
$\Psi_{Q}^{\prime}=-2\int_{0}^{\infty}\widetilde{\mathbb{E}}^{m}_{\varphi}\left[e^{-\mu e^{\gamma c}\int_{\mathbb{D}}\frac{\mathrm{d}M_{\gamma}(z)}{|z|^{\gamma Q}}}\right]\mathrm{d}m,$
[…]
All the estimates used in this work break down at $\gamma=\sqrt{2}$. It should be possible to adapt some techniques from [DKRV17, Bav19, BW18] to get more precise estimates, but the analysis looks more subtle. This is a purely probabilistic question, and we view it as an interesting problem on Gaussian multiplicative chaos.
@@
Can our techniques be adapted to general $(r,s)$? In [BW23], we relied on an induction formula to express a singular integral in terms of “lower” ones (with respect to a certain partial order on partitions). The same technique should apply to the HEM, but it will be computationally heavier, with the additional difficulty of computing the residue. In the end, we expect that the only non-zero contribution at $\alpha_{r,s}$ is given by the residue of $\int_{\mathbb{D}^{r}}\Psi_{\alpha}(w_{1},...,w_{r})\prod_{j=1}^{r}\frac{|\mathrm{d}w_{j}|^{2}}{|w_{j}|^{2s}}$. We note that the $(1,s)$-equation is the simplest and could be directly deduced from the techniques of this paper, but we prefer to treat the general case all at once.
@@
Due to a freezing phenomenon, the $(2,1)$-equation crashes down for $\gamma>\sqrt{2}$. However, even though $\alpha_{-2,1}>Q$ for $\gamma>\sqrt{2}$, the state $\Psi_{\alpha_{-2,1}}$ can be analytically continued by $R(\alpha_{-2,1})\Psi_{2Q-\alpha_{-2,1}}$ [${}^{+}$23]. Is there a probabilistic model such that the $(2,1)$-equation is non-trivial and given by this analytic continuation? Based on [AHS21, Section 3.2], one possibility would be to look at Poisson collections of quantum spheres.
@@ Acknowledgements
This project was initiated during the school “Three facets of gravity” held in spring 2023 at Humboldt University, thanks to the support of Kolleg Mathematik Physik Berlin.
@@ 2.1. Setup and background
The content of this section is extracted from [GKRV23, ${}^{+}$23], and follows closely the summary given in [BW23, Section 2].
@@ 2.1.1. Free field modules
[…]
Let $\mathbb{P}_{\mathbb{S}^{1}}$ be the law of a $\log$-correlated Gaussian field $\varphi$ on the unit circle:
$\mathbb{E}[\varphi(e^{i\theta})\varphi(e^{i\theta^{\prime}})]=\log\frac{1}{|e^{i\theta}-e^{i\theta^{\prime}}|}.$
The expansion in Fourier modes reads
$\varphi=\sum_{n\in\mathbb{Z}\setminus\{0\}}\varphi_{n}e_{n},$
[…]
@@
$\mathbb{E}[P\varphi(z)P\varphi(w)=\log\frac{1}{|1-z\bar{w}|}=:G_{\partial}(z,w),\qquad\forall z,w\in\mathbb{D}.$
@@
where $\mathrm{d}c$ is Lebesgue measure on $\mathbb{R}$. Samples of $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$ are written $c+\varphi$, with $c$ being the zero mode of the field. We define a dense subspace $\mathcal{C}$ of $\mathcal{H}$ as the subspace of functionals $F\in\mathcal{H}$ such that there exists $N\in\mathbb{N}^{*}$ and $f\in\mathcal{C}^{\infty}(\mathbb{R}\times\mathbb{C}^{N})$, such that $F(c+\varphi)=f(c,\varphi_{1},...,\varphi_{N})$ holds $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$-a.e., and $f$ and all its derivatives are compactly supported in $c$ and have at most exponential growth in the other modes. We refer to $\mathcal{C}$ as the space of test functions, and its continuous dual $\mathcal{C}^{\prime}$ as the space of tempered distributions.
Let $\alpha\in\mathbb{C}$. On $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$, we have two commuting representations of the Heisenberg algebra $(\mathbf{A}_{n},\widetilde{\mathbf{A}}_{n})_{n\in\mathbb{Z}}$ given for $n>0$ by
@@
Here, $\partial_{n}=\partial_{\varphi_{n}}$ mean (complex) derivative in direction $\varphi_{n}$ and $\partial_{-n}=\partial_{\bar{\varphi}_{n}}$. For $n\neq 0$, we have the hermiticity relations $(\mathbf{A}_{n}^{\alpha})^{*}=\mathbf{A}_{-n}^{\alpha}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. Given a partition $\mathbf{k}\in\mathcal{T}$, we set $\mathbf{A}_{-\mathbf{k}}:=\prod_{n=1}^{\infty}(\mathbf{A}_{-n}^{\alpha})^{k_{n}}$, and for $\mathbf{k},\widetilde{\mathbf{k}}\in\mathcal{T}$, we set
$\pi_{\mathbf{k},\mathbf{\widetilde{k}}}:=\mathbf{A}_{-\mathbf{k}}\widetilde{\mathbf{A}}_{-\widetilde{\mathbf{k}}}\mathds{1},$
where $\mathds{1}$ is the constant function 1. We say that $\pi_{\mathbf{k},\widetilde{\mathbf{k}}}$ has level $|\mathbf{k}|+|\widetilde{\mathbf{k}}|$. The Heisenberg representation gives $\mathcal{F}$ a structure of highest-weight Heisenberg module, i.e.
$\mathcal{F}=\mathrm{span}\,\{\pi_{\mathbf{k},\widetilde{\mathbf{k}}}|\,\mathbf{k},\widetilde{\mathbf{k}}\in\mathcal{T}\}.$
The level gives a grading
$\mathcal{F}=\oplus_{N\in\mathbb{N}}\mathcal{F}_{N},$
and $\dim\mathcal{F}_{N}=\sum_{n=0}^{N}p(n)p(N-n)$.
The Sugawara construction consists in two commuting representation $(\mathbf{L}_{n}^{0,\alpha},\widetilde{\mathbf{L}}_{n}^{0,\alpha})_{n\in\mathbb{Z}}$ of the Virasoro algebra on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. These operators are the following quadratic expression in the Heisenberg algebra ($n\neq 0$)
$\mathbf{L}_{n}^{0,\alpha}:=i(\alpha-(n+1)Q)\mathbf{A}_{n}+\sum_{m\neq\{0,n\}}\mathbf{A}_{n-m}\mathbf{A}_{m};$ $\mathbf{L}_{0}^{0,\alpha}:=\Delta_{\alpha}+2\sum_{m=1}^{\infty}\mathbf{A}_{-m}\mathbf{A}_{m};$
$\widetilde{\mathbf{L}}_{n}^{0,\alpha}:=i(\alpha-(n+1)Q)\widetilde{\mathbf{A}}_{n}+\sum_{m\neq\{0,n\}}\widetilde{\mathbf{A}}_{n-m}\widetilde{\mathbf{A}}_{m};$ $\mathbf{L}_{0}^{0,\alpha}:=\Delta_{\alpha}+2\sum_{m=1}^{\infty}\widetilde{\mathbf{A}}_{-m}\widetilde{\mathbf{A}}_{m}.$
This representation satisfies the hermiticity relations $(\mathbf{L}_{n}^{0,\alpha})^{*}=\mathbf{L}_{-n}^{0,2Q-\bar{\alpha}}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. Given a partition $\nu=(\nu_{1},...,\nu_{\ell})$, we set $\mathbf{L}_{-\nu}^{0,\alpha}=\mathbf{L}_{-\nu_{\ell}}^{0,\alpha}...\mathbf{L}_{-\nu_{1}}^{0,\alpha}$. Similar formulas and notations hold for the representation $\widetilde{\mathbf{L}}_{n}^{0,\alpha}$. The descendant state
@@
where the span is algebraic (finite linar combinations). We also define $\mathcal{V}_{\alpha}^{0,N}:=\mathcal{V}_{\alpha}^{0}\cap\mathcal{F}_{N}$. The module $\mathcal{V}_{\alpha}^{0}$ has central charge $c_{\mathrm{L}}=1+6Q^{2}$ and highest-weight $\Delta_{\alpha}=\frac{\alpha}{2}(Q-\frac{\alpha}{2})$. If $\alpha\not\in kac$, it is known that $\mathcal{V}_{\alpha}^{0}$ is irreducible and Verma (hence $\mathcal{V}_{\alpha}^{0}\simeq\mathcal{F}$) when $\alpha\not\in kac$. On the other hand, if $\alpha\in kac^{-}$, $\mathcal{V}_{2Q-\alpha}^{0}$ is Verma (hence $\mathcal{V}_{2Q-\alpha}^{0}\simeq\mathcal{F}$), and $\mathcal{V}_{\alpha}^{0}$ is the irreducible quotient of the Verma by the maximal proper submodule. The linear map
$\Phi_{\alpha}^{0}:\left\{\begin{aligned} &\mathcal{V}_{2Q-\alpha}^{0}\simeq\mathcal{F}&&\to\mathcal{V}_{\alpha}^{0}\\ &\mathbf{L}_{-\nu}^{0,2Q-\alpha}\widetilde{\mathbf{L}}_{-\nu}^{0,2Q-\alpha}\mathds{1}&&\mapsto\mathbf{L}_{-\nu}^{0,\alpha}\widetilde{\mathbf{L}}_{-\nu}^{0,\alpha}\mathds{1}\end{aligned}\right.$
@@
$\mathbf{S}^{0}_{\alpha}:=\alpha^{2}\mathbf{L}_{-2}^{0,\alpha}+(\mathbf{L}_{-1}^{0,\alpha})^{2};$
$\widetilde{\mathbf{S}}^{0}_{\alpha}:=\alpha^{2}\widetilde{\mathbf{L}}_{-2}^{0,\alpha}+(\widetilde{\mathbf{L}}_{-1}^{0,\alpha})^{2}.$
@@
$\mathbf{S}^{0}_{\alpha}=2\alpha(\alpha^{2}+\alpha Q+1)\varphi_{2}=2\alpha(\alpha-\alpha_{1,2})(\alpha-\alpha_{2,1})\varphi_{2};$
@@
$\widetilde{\mathbf{S}}^{0}_{\alpha}\mathbf{S}^{0,\alpha}=\alpha^{2}(\alpha-\alpha_{1,2})^{2}(\alpha-\alpha_{2,1})^{2}(4|\varphi_{2}|^{2}-1).$

### v2

@@ 1.2.2. Boundary HEMs
Following the same mechanism as in the bulk case, Theorem 1.2 should lead to PDEs for correlation functions on arbitrary Riemann surfaces with boundaries and punctures. This will follow from a series of works on boundary Liouville theory [GRW24b, 33]. As mentioned in Section 1.1, [ANG23] proves the BPZ equations at the level of correlation functions of Liouville CFT in the disc, which corresponds to the cases where the right-hand-sides in Theorem 1.2 vanish. The proof of [ANG23] is quite different from ours since it uses the mating-of-trees theory [DMS21]: it relies on the clever observation that the BPZ differential operator is the generator of a certain Schramm–Loewner evolution. Using this observation, the BPZ equation is a consequence of the fact that the correlation function evolves as a martingale under the dynamics generated by this operator. In this context, the quantities $\cos(\pi\frac{\gamma^{2}}{4})$ and $\sin(\pi\frac{\gamma^{2}}{4})$ appearing in the right-hand-side of the $(2,1)$-HEM have a special meaning: they are respectively the correlation and the speed of the Brownian motions corresponding to the evolution of the two quantum lengths (parametrised by quantum area). In our proof, these quantities are related to residues of certain Selberg integrals on the degenerate weights (Lemma 3.6).
Finally, [CER25] also works at the level of correlation functions, but proves all HEMs in a way that matches Theorem (1.2). The method is closer to earlier works in the field, using the conformal Ward identities. Cerclé introduces a “descendant field at level 2” which can be thought of as the descendant state $\Phi_{\alpha}(\mathbf{S}^{0}_{2Q-\alpha}\mathds{1})$ from this paper. The insertion of this state combined with Ward identities leads to the BPZ equation. In our approach, the Ward identities will be replaced with the conformal blocks machinery that will be developed thanks to the boundary version of the conformal bootstrap. Indeed, this machinery gives a way to streamline the Ward identities without computations and independently of the geometric setup.
1.3. Outline and open questions
The strategy for the proof of Theorems 1.1 and 1.2 is the following.
@@ 1.3. Outline and open questions
    We express the singular state as certain singular integral of the correlation function (Sections 2.2 and 3.2). These expressions are reminiscent of those found in [BW23], namely they look like primary fields with additional $\gamma$-insertions integrated over the disc (see (2.5) and (3.3)).
@@
    Contrary to [BW23], these integrals have a simple pole at $\alpha_{1,2}$, $\alpha_{2,1}$. The presence of this pole is the crux of the HEM: the RHS of the HEMs are given by the residues of these integrals. To compute this residue, we rely on fusion estimates for correlation functions (Propositions 2.6 and 3.7), and exact formulas for Dotsenko–Fateev/Selberg integrals.
@@
    The theory developed in [GKR+24a, BGK+24a] tells us that the Poisson operator is analytic away from the Kac table. Since $\alpha_{1,2}<\alpha_{2,1}$, the pole at $\alpha_{1,2}$ is easier to study. To study the pole at $\alpha_{2,1}$, one needs to find a probabilistic expression for the analytic continuation beyond the pole at $\alpha_{1,2}$. This done in Propositions 2.7 and 3.8, using a method similar to the one in [BW23, Section 3.3].
@@
The RHS of (1.3) has a limit as $\gamma\to\sqrt{2}$ from below, but this limit is deceptively simple: the constant prefactor has a pole (as a function of $\gamma$), which is compensated by the zero of the primary state. Thus, the limit involves the derivative state $\Psi_{Q}^{\prime}=\underset{\alpha\nearrow Q}{\lim}\,\frac{\Psi_{\alpha}}{\alpha-Q}$ corresponding to the “$Q$-vertex operator” studied in [DKR+17]. The estimates used in this paper are not sufficient to prove the case $\gamma=\sqrt{2}$, but it should be possible to adapt some techniques from [DKR+17, BAV19, BW25] to get more precise estimates (although the analysis looks more subtle). We view it as an interesting question on Gaussian multiplicative chaos.
@@
Can our techniques be adapted to general $(r,s)$? In [BW23], we relied on an induction formula to express a singular integral in terms of “lower” ones (with respect to a certain partial order on Young diagrams). The same technique should apply to the HEM, but it will be computationally heavier, with the additional difficulty of computing the residue. In the end, we expect that the only non-zero contribution at $\alpha_{r,s}$ is given by the residue of $\int_{\mathbb{D}^{r}}\Psi_{\alpha}(w_{1},...,w_{r})\prod_{j=1}^{r}\frac{|\mathrm{d}w_{j}|^{2}}{|w_{j}|^{2s}}$. We note that the $(1,s)$-equation is the simplest and could be directly deduced from the techniques of this paper, but we prefer to treat the general case all at once.
@@
Due to a freezing phenomenon, the $(2,1)$-equation crashes down for $\gamma>\sqrt{2}$. However, even though $\alpha_{-2,1}>Q$ for $\gamma>\sqrt{2}$, the state $\Psi_{\alpha_{-2,1}}$ can be analytically continued by $R(\alpha_{-2,1})\Psi_{2Q-\alpha_{-2,1}}$ [BGK+24a]. Is there a probabilistic model such that the $(2,1)$-equation is non-trivial and given by this analytic continuation? Based on [AHS24, Section 3.2], one possibility would be to look at Poisson collections of quantum spheres.
@@ Acknowledgements
This project was initiated during the school “Three facets of gravity” held in spring 2023 at Humboldt University, thanks to the support of Kolleg Mathematik Physik Berlin. We thank the anonymous referees for their detailed feedback, which led to many improvements of the manuscript. G.B. acknowledges support from the ANR 21-CE40-0003. B.W. was supported by National Key R&D Program of China (No. 2023YFA1010700).
@@ 2.1. Setup and background
The content of this section is extracted from [GKR+24a, BGK+24a], and follows closely the summary given in [BW23, Section 2.2-3]. We also refer to [BW23, Appendix B] for some elementary background on the Virasoro algebra.
@@ 2.1.1. Free field modules
Let $\mathcal{F}:=\mathbb{C}[(\varphi_{n},\bar{\varphi}_{n})_{n\;\geqslant\;1}]$ be the space of polynomials in countably many complex variables $\varphi_{n}$ and their complex conjugates $\bar{\varphi}_{n}$. The constant function $\mathds{1}$ is the vacuum vector. A Young diagram of length $\ell\in\mathbb{N}$ is a non-increasing sequence of $\ell$ positive integers $\nu=(\nu_{1},...\nu_{\ell})$. We denote by $|\nu|=\sum_{l=1}^{\ell}\nu_{l}$ and call it the level of $\nu$. We denote the set of all Young diagrams (resp. Young diagrams of level $N$) by $\mathcal{T}$ (resp. $\mathcal{T}_{N}$) and set $p(N):=\#\mathcal{T}_{N}$. By convention, $p(0)=1$.
Let $\mathbb{P}_{\mathbb{S}^{1}}$ be the law of the $\mathbb{R}$-valued, $\log$-correlated Gaussian field $\varphi$ on the unit circle:
$\mathbb{E}[\varphi(e^{i\theta})\varphi(e^{i\theta^{\prime}})]=\log\frac{1}{|e^{i\theta}-e^{i\theta^{\prime}}|},\qquad\theta,\theta^{\prime}\in\mathbb{R}.$
Due to the logarithmic blow-up on the diagonal, this field only exists (almost surely) as a random distribution in the Sobolev space $H^{-s}(\mathbb{S}^{1})$ for arbitrary small $s>0$. The expansion in Fourier modes reads
$\varphi=\sum_{n\in\mathbb{Z}\setminus\{0\}}\varphi_{n}e_{n},$ (2.1)
where $e_{n}(e^{i\theta})=e^{ni\theta}$ is the standard basis. We have $\varphi_{-n}=\bar{\varphi}_{n}$ for all $n\in\mathbb{Z}\setminus\{0\}$. Under $\mathbb{P}_{\mathbb{S}^{1}}$, the sequence $(\varphi_{n})_{n\;\geqslant\;1}$ is made of independent complex Gaussians $\mathcal{N}_{\mathbb{C}}(0,\frac{1}{2n})$. The harmonic extension of $\varphi$ to the unit disc is
@@
$\mathbb{E}[P\varphi(z)P\varphi(w)]=\log\frac{1}{|1-z\bar{w}|}=:G_{\partial}(z,w),\qquad\forall z,w\in\mathbb{D}.$
@@
where $\mathrm{d}c$ is Lebesgue measure on $\mathbb{R}$. Samples of $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$ are written $c+\varphi$, with $c$ being the zero mode of the field. Following [GKR+24a, Section 5.2], we define a dense subspace $\mathcal{C}$ of $\mathcal{H}$ as the subspace of functionals $F\in\mathcal{H}$ for which there exist $N\in\mathbb{N}^{*}$ and $f\in\mathcal{C}^{\infty}(\mathbb{R}\times\mathbb{C}^{N})$ such that $F(c+\varphi)=f(c,\varphi_{1},...,\varphi_{N})$ holds $\mathrm{d}c\otimes\mathbb{P}_{\mathbb{S}^{1}}$-almost everywhere. Moreover, we require that $f$ and all its derivatives decay faster than any polynomial in the variable $e^{-|c|}$, and have at most polynomial growth in the other variables. We refer to $\mathcal{C}$ as the space of test functions; this space comes with a natural Fréchet topology, and we denote by $\mathcal{C}^{\prime}$ the space of continuous linear forms on $\mathcal{C}$.
Let $\alpha\in\mathbb{C}$. We have two commuting representations of the Heisenberg algebra $(\mathbf{A}_{n},\widetilde{\mathbf{A}}_{n})_{n\in\mathbb{Z}}$ acting as densely defined operators on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$, with an expression given for $n>0$ by
@@
Here, $\partial_{n}=\partial_{\varphi_{n}}$ mean (complex) derivative in the direction $\varphi_{n}$ and $\partial_{-n}=\partial_{\bar{\varphi}_{n}}$. For $n\neq 0$, we have the hermiticity relations $(\mathbf{A}_{n}^{\alpha})^{*}=\mathbf{A}_{-n}^{\alpha}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$.
The Sugawara construction consists of two commuting representations $(\mathbf{L}_{n}^{0,\alpha},\widetilde{\mathbf{L}}_{n}^{0,\alpha})_{n\in\mathbb{Z}}$ of the Virasoro algebra on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. These operators are the following quadratic expressions in the Heisenberg algebra, for $n\in\mathbb{Z}\setminus\{0\}$ (recall also (1.1)):
$\mathbf{L}_{n}^{0,\alpha}=i(\alpha-(n+1)Q)\mathbf{A}_{n}+\sum_{m\neq\{0,n\}}\mathbf{A}_{n-m}\mathbf{A}_{m};$ $\mathbf{L}_{0}^{0,\alpha}=\Delta_{\alpha}+2\sum_{m=1}^{\infty}\mathbf{A}_{-m}\mathbf{A}_{m};$ (2.2)
$\widetilde{\mathbf{L}}_{n}^{0,\alpha}=i(\alpha-(n+1)Q)\widetilde{\mathbf{A}}_{n}+\sum_{m\neq\{0,n\}}\widetilde{\mathbf{A}}_{n-m}\widetilde{\mathbf{A}}_{m};$ $\mathbf{L}_{0}^{0,\alpha}=\Delta_{\alpha}+2\sum_{m=1}^{\infty}\widetilde{\mathbf{A}}_{-m}\widetilde{\mathbf{A}}_{m}.$
This representation satisfies the hermiticity relations $(\mathbf{L}_{n}^{0,\alpha})^{*}=\mathbf{L}_{-n}^{0,2Q-\bar{\alpha}}$ on $L^{2}(\mathbb{P}_{\mathbb{S}^{1}})$. Given a Young diagram $\nu=(\nu_{1},...,\nu_{\ell})$, we set $\mathbf{L}_{-\nu}^{0,\alpha}=\mathbf{L}_{-\nu_{\ell}}^{0,\alpha}...\mathbf{L}_{-\nu_{1}}^{0,\alpha}$. Similar formulas and notations hold for the representation $\widetilde{\mathbf{L}}_{n}^{0,\alpha}$. The descendant state
@@
where the span is algebraic (finite linear combinations). We also define $\mathcal{V}_{\alpha}^{0,N}:=\mathcal{V}_{\alpha}^{0}\cap\mathcal{F}_{N}$. The module $\mathcal{V}_{\alpha}^{0}$ has central charge $c_{\mathrm{L}}=1+6Q^{2}$ and highest-weight $\Delta_{\alpha}$. If $\alpha\not\in kac$, it is known that $\mathcal{V}_{\alpha}^{0}$ is irreducible and Verma (hence $\mathcal{V}_{\alpha}^{0}\simeq\mathcal{F}$ as vector spaces) when $\alpha\not\in kac$ [KRR13, Lecture 8] (see also [BW23, Appendix B]). On the other hand, if $\alpha\in kac^{-}$, $\mathcal{V}_{2Q-\alpha}^{0}$ is Verma (hence $\mathcal{V}_{2Q-\alpha}^{0}\simeq\mathcal{F}$), and $\mathcal{V}_{\alpha}^{0}$ is the irreducible quotient of the Verma by the maximal proper submodule [FRE92] (see also [BW23, Section 2.3]). The linear map
$\Phi_{\alpha}^{0}:\left\{\begin{aligned} &\mathcal{V}_{2Q-\alpha}^{0}\simeq\mathcal{F}&&\to\mathcal{V}_{\alpha}^{0}\\ &\mathbf{L}_{-\nu}^{0,2Q-\alpha}\widetilde{\mathbf{L}}_{-\nu}^{0,2Q-\alpha}\mathds{1}&&\mapsto\mathbf{L}_{-\nu}^{0,\alpha}\widetilde{\mathbf{L}}_{-\nu}^{0,\alpha}\mathds{1}\end{aligned}\right$. (2.3)
@@
$\mathbf{S}^{0}_{\alpha}:=\alpha^{2}\mathbf{L}_{-2}^{0,\alpha}+(\mathbf{L}_{-1}^{0,\alpha})^{2}\qquad\text{and}\qquad\widetilde{\mathbf{S}}^{0}_{\alpha}:=\alpha^{2}\widetilde{\mathbf{L}}_{-2}^{0,\alpha}+(\widetilde{\mathbf{L}}_{-1}^{0,\alpha})^{2}.$
@@
$\mathbf{S}^{0}_{\alpha}\mathds{1}=2\alpha(\alpha^{2}+\alpha Q+1)\varphi_{2}=2\alpha(\alpha-\alpha_{1,2})(\alpha-\alpha_{2,1})\varphi_{2};$
@@
$\widetilde{\mathbf{S}}^{0}_{\alpha}\mathbf{S}^{0}_{\alpha}\mathds{1}=\alpha^{2}(\alpha-\alpha_{1,2})^{2}(\alpha-\alpha_{2,1})^{2}(4|\varphi_{2}|^{2}-1).$